The perimeter of a rectangle is 60 inches. The ratio of the length to width is 3:2. What is the length and width?
Answer by ad_alta(240)

Let 'w' be the width and 'l' be the length. Then the perimeter is the sum of the sides of the rectangle. Thus, 2w+2l=60. Since l/w=3/2, 2l=3w. We can substitute into the first equation and get 2w+2(3/2)w=60. Then 5w=60 and w=12. Since l=(3/2)w, l=18.